OpenStudy (spring98):

Multiply (2.0 • 10^–4) • (3.1 • 10^–20). Express the answer in scientific notation.

OpenStudy (spring98):

I got 6.2 *10^80. AM I right @CShrix ?

OpenStudy (jiteshmeghwal9):

\[(2 \times 10^{-4})(3.1 \times 10^{-20})\]\[6.2 \times 10^{-4-20}=6.2 \times 10^{-24}\]

OpenStudy (anonymous):

Don't forget, whenever you multiply exponents with the same base, you add them!! \[\huge x^2 \cdot x^4=x^6\]So\[\huge 10^{-20} \cdot 10^{-4}=10^{-24}\]

OpenStudy (spring98):

thank you very much!!! I wish I knew how to use the equation thing better on os:{

OpenStudy (anonymous):

Haha, no worries :)